find The mass of CO2, that must be mixed with 20 g of oxygen such that 27 ml of a sample of the resulting mixture would contain equal number of molecules of each gas

Grade:11

2 Answers

Gopal
6 years ago
Equal moles contain equal number of particles 
number of moles of oxygen n =\tfrac{w}{M} = 20/32 =0.625
therefore mass of carbondioxide mixed should be =0.625*44 =27.5 g
Now ,total volume contain equal number of molecules of oxygen and carbon dioxide.
So, any fraction of volume also contain equal number of molecules.
Shalini Ravikant Kshirsagar
4 years ago
CO2 be x gm
no.of moles of O2 =given mass ÷molar mass
                      =20/32
                       =0.625 mol
Now,for for 0.625 moles of CO2
   For 1mol ,44gmso, 
So,for 0.625 mol ,x gm
         X=44*0.625/1
         X=27.5 gm